OpenStudy (anonymous):

Help!!! Linear Equations {2x+3y=13 {15x-3y+21

OpenStudy (anonymous):

what are you trying to find? the solution set?

OpenStudy (anonymous):

x=-8/17 y=79/17

OpenStudy (anonymous):

ugly solution set

OpenStudy (anonymous):

The instructions say: Use the Elimination Method to solve each system of linear equations.

OpenStudy (chriss):

add the bottom to the top and you get 17x=34 solve for x and you get 2 then sub 2 into one of the original equations to get y which will get you 3 so x=2 and y=3

OpenStudy (anonymous):

it wouldn't be that hard to solve by hand since your y's are already set up

OpenStudy (anonymous):

15*2+-3*3=-21? unless that plus sign is supposed to be an equal sign

myininaya (myininaya):

+== i believe

OpenStudy (anonymous):

depending on whether that is a plus or an equal sign lol

OpenStudy (chriss):

I didn't see that, but I do think the + is suposed to be an =

OpenStudy (anonymous):

wow party on this thread?

myininaya (myininaya):

we can use the operations anyway we want as long as we define those operations

OpenStudy (chriss):

If it's a system of equations set up to be an easy elimination problem then =21 makes sense :P

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

it'd be easy still if it was a plus sign also it'd just be 15x-3y=-21

myininaya (myininaya):

richellea which is it lol you know you don't have an equal sign in your second line right?

OpenStudy (anonymous):

either way we are both correct he can distinguish between the two answers XD

OpenStudy (anonymous):

Yes there is an equal sign.